a recipe says to mix 480 ml of milk with 100ml of water how much water should i use if i only have 360 ml of milk
Answer by checkley77(12844)

480/100=360/X
480X=100*360
480X=36000
X=36000/480
X=75 ML. OF WATER IS USED.